# Expedited Noise Spectroscopy of Transmon Qubits

### v1

@@ Abstract
There has been tremendous progress in the physical realization of quantum computing hardware in recent times, bringing us closer than ever before to realizing the promise of quantum computing. However, noise continues to pose a crucial challenge when it comes to scaling up present day quantum processors. While decoherence limits the qubits’ ability to store information for long periods in the presence of uncontrollable noise sources, the erroneous implementation of control methods for state preparation and measurements leads to faulty implementations of quantum circuits. Conventional noise spectroscopy protocols can characterize and model environmental noise but are usually resource intensive and lengthy. Moreover, the underlying noise can vary in nature over time, making noise profile extraction futile as this new information cannot be harnessed to improve quantum error correction or dynamical decoupling protocols. In this work, we address this challenge using a machine learning-based methodology to quickly extract noise spectra of multiple qubits and demonstrate a possible noise mitigation strategy. The procedure involves implementing undemanding dynamical decoupling sequences to record coherence decays of the investigated qubits and then predict the underlying noise spectra with the help of a convolution neural network pre-trained on a synthetic dataset. While our protocol is virtually hardware-agnostic, we validate its effectiveness using superconducting qubits available on the IBM Quantum platform. We further use these rapidly obtained, yet accurate, noise spectra to design bespoke dynamic decoupling sequences and perform time-dependent noise spectroscopy.
@@ I INTRODUCTION
Decoherence Zurek (2003); Schlosshauer (2005); Ithier et al. (2005) refers to the loss of coherence in a quantum system due to unavoidable interactions with its environment. Decoherence remains the Achilles heel in developing practical quantum technologies such as quantum computation Putterman et al. (2024); Dalzell et al. (2023), and quantum sensing Dolde et al. (2011); Degen et al. (2017). Understanding and controlling noise leading to decoherence in quantum systems in therefore an important task in the context of building robust and scalable quantum processors. To this end, one needs to understand the nature of this noise, identify its possible sources, and efficiently curb it with the available tools.
@@
Characterizing and modeling noise using conventional spectroscopic methods, requires abundant resources. The way to extract a spectrum of noise that affects a system is based on deconvolution of the spectral overlap of the noise power spectral density function and the corresponding quantum control protocol used to probe the system Szankowski et al. (2017). This control method can also be a dynamical error suppression protocol, such as dynamical decoupling (DD) Cywiński et al. (2008); Hernández-Gómez et al. (2018); Bar-Gill et al. (2012); Medford et al. (2012); Khodjasteh and Lidar (2005); Carr and Purcell (1954); Hahn (1950); Uhrig (2007), which is a general method to preserve spin coherence in the presence of noise. The control acts as a filter, allowing one to probe a specific region of the broad frequency spectrum affecting the quantum system. This can be modeled analytically with the help of the filter function formalism Szankowski et al. (2017); Uys et al. (2009). Conversely, understanding the spectrum of the noise acting on a qubit crucially underpins the optimization of DD protocols that can mitigate such noise. In practice, the noise spectrum varies significantly across different qubits in non-trivial ways that are hard to predict or accurately extract from most common measurements. As a result, it is difficult to predict a priori which of the several possible DD protocols would provide optimal suppression of decoherence. Indeed, one could imagine constructing a decoupling protocol customized for a particular qubit, but this is possible only if the knowledge of the actual qubit noise spectrum is available with sufficient accuracy. Furthermore, the design and implementation of efficient quantum error correction (QEC) Terhal (2015); Jayashankar and Mandayam (2023); Chou et al. (2024) protocols and fault-tolerant quantum circuits Heußen et al. (2024), often requires an understanding of the specific noise affecting the qubits; rapid extraction of noise spectra thus becomes an imperative tool for both error mitigation and error correction.
@@
Extracting noise spectra rapidly for qubits presents several challenges. Conventional long-duration noise spectroscopy protocols require extended measurement times and repeated implementation Álvarez and Suter (2011). Thus, for time-dependent noise, the measured noise spectrum becomes a distorted representation of the actual noise affecting the qubit. Furthermore, other complex pulse sequences with minimal spectral leakage can allow noise extraction with relatively fewer experimental runs Norris et al. (2018); however, these strategies require unique resources and usually have reduced sensitivity. Here, we address the challenge of rapid noise spectroscopy with the help of deep learning algorithms. Owing to their success in diverse domains, there has been a thrust to utilize the power of artificial neural networks to characterize and correct the deleterious environmental effects Wise et al. (2021); Meneses et al. (2022); Papič and de Vega (2022); Canonici et al. (2023); Lu et al. (2025). In this study, we extend the neural network-based methodology Wise et al. (2021); Meneses et al. (2022) to rapidly predict the time-varying noise spectral densities associated with the state-of-the-art transmon qubits. Moreover, we propose a proof-of-principle method to mitigate the environmental noise by constructing customized dynamical decoupling sequences.
@@
The rest of the paper is organized as follows. In Sec. II, we discuss the mathematical framework for modeling noise in quantum systems and highlight the importance of accurately discerning the environmental noise. We also recall the link between the extracted noise spectrum and the error rates corresponding to various noise sources Bylander et al. (2011). Sec. III details our methodology. Therein, we demonstrate how a deep neural network can be trained to extract the noise spectrum from typical time dynamics measurements such as Hahn echo and Carr-Purcell-Meiboom-Gill (CPMG)-32. We then provide an overview of the experimental data processing and DD optimization procedure. In Sec. IV, we benchmark our model’s effectiveness in extracting and mitigating the noise for the IBM Quantum superconducting quantum architecture. Finally, in Sec. V, we conclude with critical discussion and propose future directions.
@@ II Significance of the ‘Decoherence Functional’
Sources Arises from Type of coupling with qubit S($\omega$)
Charge Noise Charge fluctuators present in the Dominant: Transverse ($T_{1}$) At low frequency $\propto 1/\omega$
defects or charge traps that resi- Sometimes low frequency ($T_{\phi}$) At high frequency $\propto\omega$
de in interfacial dielectrics, the
junction tunnel barrier, and in
the substrate itself.
Magnetic Flux Noise Stochastic flipping of qubits that Transverse ($T_{1}$)
reside on the surfaces of the supe- Longitudinal ($T_{\phi}$)
rconductors, resulting in random $\propto 1/\omega$
fluctuations of effective-$\vec{B}$ that
biases flux-tunable qubits.
Photon Number Fluctuations In resonator, residual microwave Longitudinal ($T_{\phi}$) $\propto\frac{k}{\omega^{2}+k^{2}}$
fields in the cavity have photon $k:$ resonator decay rate
number fluctuations.
Quasiparticles Unpaired electrons tunneling th- Transverse ($T_{1}$) $\propto\frac{N_{qp}\tau_{qp}}{\omega^{2}+\tau_{qp}^{2}}$
rough a qubit junction Longitudinal ($T_{\phi}$) $\tau_{qp}:$ quasiparticle lifetime
Local two level system (TLS) Electric dipole moment resonate- Transverse ($T_{1}$) Low freq.: White noise, $1/\omega$
ly absorb energy from the oscill- Longitudinal ($T_{\phi}$) High freq.: Lorentzian type
ating $\vec{E}$ of the qubit mode, and
efficiently dissipate it into the
phonon or quasiparticle bath.
Table 1: Summary of common noise sources in transmon systems Krantz et al. (2019), their frequency dependencies, and coupling mechanisms with qubits.
@@
[…]
Here, the $T_{1}$ time corresponds to the energy relaxation process, while $T_{\phi}$ represents pure dephasing and $\Gamma_{\phi}=1/T_{\phi}$ is the pure dephasing rate. The total decoherence time $T_{2}$ reflects the loss of coherence caused by both energy relaxation and pure dephasing. The noisy density matrix $\rho_{\rm noisy}$ for the qubit after the impact of noise can then be written as Nielsen and Chuang (2011),
$\rho=$ $\begin{pmatrix}|\alpha|^{2}&\alpha\beta^{*}\\ \alpha^{*}\beta&|\beta|^{2}\end{pmatrix}\xrightarrow{\rm noise}$
$\rho_{\rm noisy}=\begin{pmatrix}1+(|\alpha|^{2}-1)e^{-\Gamma_{1}t}&\alpha\beta^{*}e^{-\Gamma_{2}t}\\ \alpha^{*}\beta e^{-\Gamma_{2}t}&|\beta|^{2}e^{-\Gamma_{1}t}\end{pmatrix}$ (1)
@@
For $1/f$ noise spectra, the phase decay function is itself a Gaussian $\exp[-(t/T_{\phi,G})^{2}]$ (stretched exponential decay) Krantz et al. (2019). Furthermore, this function is separable from the $T_{1}$-type exponential decay because the $T_{1}$-noise remains regular at the qubit frequency. Therefore, the modified density matrix after the action of noise is now given by Krantz et al. (2019),
@@
Here, $C(t)$ denotes the coherence curve and $F(\omega t)$ is the filter function associated with a given pulse sequence. In addition to $1/f$-type dephasing mechanisms, there are also “white” pure dephasing mechanisms, which give rise to an exponential decay function for the dephasing component of $T_{2}$. The noise spectrum $S(\omega)$ thus exclusively characterizes the dephasing mechanism.
It is important to note here that the noisy density matrix in Eq. (II) is equivalent to what one may obtain using the Kraus operator-sum description of noise in quantum systems Nielsen and Chuang (2011). The Kraus operators describe the decoherence resulting from the interaction between the system and its environment over a period of time. The form of the Kraus operators takes into account the decoherence functional, as explained in Appendix A. The density matrix in Eq. (II) is thus identical to the density matrix in Eq. (A) obtained due to the combined effect of the dephasing and amplitude-damping channels on the initial density matrix, assuming that the dephasing probability $p(t)$ is simply given by $p=\frac{1}{2}(1+e^{-\Gamma_{\phi}t})$.
@@
In this case, the noisy density matrix in Eq. (A) gets modified to Eq. (2). This elucidates the connection between the two different approaches by demonstrating how the decoherence functional can be utilized to derive the explicit probabilities within the Kraus operator framework.
This connection makes a more precise simulation of noisy dynamics possible, bridging the gap between theoretical descriptions and practical implementations of quantum systems under the influence of noise. Therefore, critical information on qubit dynamics can be derived from the noise spectral density, $S(\omega)$, which characterizes the environment. Extracting $S(\omega)$ from the measured signal, however, is not straightforward, as the process typically involves a deconvolution prone to error. Recent work Meneses et al. (2022) suggests that this problem can be largely mitigated through the use of deep feedforward neural networks as explained in the subsequent section.
@@ III Noise Spectroscopy & Mitigation Methodology
Figure 1: (a) Randomly chosen 5 test input decoherence curves as functions of evolution time; the dashed lines denote the curves evaluated from the predicted $S(\omega)$. (b) A schematic diagram of the neural network developed in this work. (c) The numerically generated (solid lines) and the predicted (dashed line) noise spectra associated with the input decoherence curves. The inset shows a histogram of the estimation errors.
The proposed methodology is designed specifically to extract the noise spectrum $S(\omega)$ associated with $T_{\phi}$ relaxation noise. Experimentally, the $T_{\phi}$ decay function is isolated by removing the $T_{1}$ contribution from the overall $T_{2}$ decay. While $T_{1}$ relaxation can be addressed primarily through improved qubit fabrication and quantum error correction, our focus here is on estimating and mitigating dephasing noise.
@@
    Training Data Generation: Based on the previous work Krantz et al. (2019), we assumed that the noise spectra $S(\omega)$ follow some complex yet well-defined functional form. Particularly, we considered that white noise dominates at relatively low frequencies, $1/\omega$-type noise becomes prominent at intermediate frequencies, and $k/(k^{2}+\omega^{2})$-type noise takes over at relatively high frequencies. Thus, we generated tens of thousands of noise spectral densities $S(\omega)$ by varying noise amplitudes and frequency cutoffs associated with the aforementioned noise types. Additionally, the stitched noise spectra were smoothed to avoid unrealistic discontinuities. We then used Eq. (3) to evaluate the corresponding decoherence curves $C(t)$ by convoluting the generated noise spectra with the CPMG-32 filter function. Finally, we addded random noise to the coherence curves $C(t)$ to simulate experimental noise. More details on data generation can be found in the Appendix B.
@@
    Network Construction and Training: Using the numerically generated noise spectral densities and their associated coherence curves $C(t)$, we trained a convolution neural network Meneses et al. (2022) to identify the noise spectral density $S(\omega)$ based on a single decoherence curve $C(t)$ provided at the network input. A detailed description is provided in Appendix B .
@@
    Experimental Data Acquisition: For experimental validation, we performed measurements on IBM’s superconducting quantum processors using the Qiskit package. After determining the optimal amplitude for the square-shaped $\pi$-pulse, we performed $T_{1}$ and $T_{2}$ decay experiments using a CPMG-32 pulse protocol. The choice of CPMG-32 sequence over conventional Hahn-echo protocol was made to enable probing of the complex high-frequency region of the noise spectrum, which lies beyond the white noise-dominated low-frequency region. Using the $T_{1}$ and $T_{2}$ data we generated the pure dephasing $T_{\phi}$ decay as per Eq. (5), where $P_{0}$ and $P_{1}$ are the probabilities of $|0\rangle$ (for $T_{2}$) and $|1\rangle$ (for $T_{1}$):

### v2

@@ Abstract
There has been tremendous progress in the physical implementation of quantum protocols in recent times, bringing us closer than ever to realizing the promise of quantum computing. However, environmental noise continues to pose a crucial challenge to scaling up present-day quantum processors. In the presence of uncontrollable noise sources, decoherence limits the qubits’ ability to store information for long periods. Conventional noise spectroscopy protocols can characterize and model environmental noise but are usually resource-intensive and lengthy. Moreover, the noise can vary in time, making the slow extraction futile as the profile cannot be harnessed to perform error mitigation or correction. Here, this challenge is addressed using a machine learning-based methodology that outputs the noise associated with transmon qubits with minimal absolute error. The procedure involves implementing undemanding dynamical decoupling sequences to record coherence decays of the qubits and then predicting the underlying noise spectra with the help of a convolutional neural network pre-trained on a synthetic dataset. While the protocol is virtually hardware-agnostic, its effectiveness is validated using superconducting qubits available on the IBM Quantum platform. These rapidly obtained, yet accurate, noise spectra are further used to design bespoke dynamic decoupling sequences and perform time-dependent noise spectroscopy
@@ I INTRODUCTION
Decoherence Zurek (2003); Schlosshauer (2005); Ithier et al. (2005) refers to the loss of coherence in a quantum system due to unavoidable interactions with its environment. It remains the Achilles heel in developing practical quantum technologies such as quantum computation Putterman et al. (2025); Dalzell et al. (2023), and quantum sensing Dolde et al. (2011); Degen et al. (2017). Understanding and controlling noise leading to decoherence in quantum systems is therefore an important task in the context of building robust and scalable quantum processors. To this end, one needs to understand the nature of this noise, identify its possible sources, and efficiently curb it with the available tools.
@@
Characterizing and modeling noise using conventional spectroscopic methods, requires abundant resources. The way to extract a spectrum of noise that affects a system is based on deconvolution of the spectral overlap of the noise power spectral density function and the corresponding quantum control protocol used to probe the system Szankowski et al. (2017). This control method can also be a dynamical error suppression protocol, such as dynamical decoupling (DD) Hahn (1950); Carr and Purcell (1954); Meiboom and Gill (1958); Khodjasteh and Lidar (2005); Uhrig (2007); Medford et al. (2012); Bar-Gill et al. (2012); Hernández-Gómez et al. (2018), which is a general method to preserve spin coherence in the presence of noise. The control acts as a filter, allowing one to probe a specific region of the broad frequency spectrum affecting the quantum system. This can be modeled analytically with the help of the filter function formalism Szankowski et al. (2017); Uys et al. (2009). Conversely, understanding the spectrum of the noise acting on a qubit crucially underpins the optimization of DD protocols that can mitigate such noise. In reality, the noise spectrum varies substatially across different qubits in complex and unpredictable ways that are difficult to accurately extract from standard measurements. Therefore, it is hard to determine in advance, which of the several possible DD protocols would offer optimal suppression of decoherence. One could imagine developing a sequence tailored to a particular qubit, but this is achievable only if detailed and accurate understanding of the underlying qubit noise spectrum can be obtained. Furthermore, the design and implementation of efficient quantum error correction (QEC) Terhal (2015) protocols and fault-tolerant quantum circuits Heußen et al. (2024), can improve significantly with an understanding of the specific noise affecting the qubits Leung et al. (1997); Tuckett et al. (2019); Chou et al. (2024); Biswas et al. (2025); Dutta et al. (2024); rapid extraction of noise spectra thus becomes an imperative tool for both error mitigation and error correction.
@@
Extracting noise spectra rapidly for qubits presents several challenges. While standard quantum noise spectroscopy (QNS) protocols based on Pauli measurements have enabled significant progress in characterizing both single- and multi-qubit noise correlations, Norris et al. (2016); Paz-Silva et al. (2017); Sung and et al (2019); Sung et al. (2021); von Lüpke et al. (2020); Álvarez and Suter (2011) these methods are generally time-intensive and require complex post-processing. Additionally, conventional QNS protocols are highly effective for stationary noise environments, their extended measurement times can lead to a distorted representation of the actual noise spectrum in the presence of time-dependent noise. Furthermore, other complex pulse sequences with minimal spectral leakage can allow noise extraction with relatively fewer experimental runs; Norris et al. (2018) however, these strategies require unique resources and usually have reduced sensitivity. To mitigate these challenges, our approach leverages standard experimental protocols, such as Carr-Purcell-Meiboom-Gil (CPMG) sequences Hahn (1950); Carr and Purcell (1954); Meiboom and Gill (1958) and utilizes a convolutional neural network to rapidly and robustly infer the underlying noise spectrum from measured coherence decay curves. Owing to their success in diverse domains, there has been a thrust to utilize the power of artificial neural networks to characterize and correct the deleterious environmental effects. Wise et al. (2021); Meneses et al. (2022); Papič and de Vega (2022); Canonici et al. (2023); Lu et al. (2025) We note that Youssry et al.Youssry et al. (2020) have devised a machine learning based ‘graybox’ approach that focuses on the characterization of the dynamics of a two-level open quantum system. However, the process of generating training data and implementing the approach on realistic hardware is non-trivial. In this study, we extend a distinct neural network driven methodology Wise et al. (2021); Meneses et al. (2022) to rapidly predict the time-varying noise spectral densities associated with the state-of-the-art transmon qubits. The technique allows us to potentially capture transient features and minimize temporal averaging, thereby providing a more accurate reflection of the instantaneous noise affecting the qubit. This is specifically useful for fast and scalable device benchmarking, complementing existing QNS techniques. Moreover, we propose a proof-of-principle method to mitigate the environmental noise by constructing customized dynamical decoupling sequences.
@@
The rest of the paper is organized as follows. In Section II, we discuss the mathematical framework for modeling noise in quantum systems and highlight the importance of accurately discerning the environmental noise. We also recall the link between the extracted noise spectrum and the error rates corresponding to various noise sources Bylander et al. (2011). Section III details our methodology. Therein, we demonstrate how a deep neural network can be trained to extract the noise spectrum from typical time dynamics measurements such as Hahn echo and CPMG. We then provide an overview of the experimental data processing and DD optimization procedure. In Section IV, we benchmark our model’s effectiveness in extracting and mitigating the noise for the IBM Quantum superconducting quantum architecture. Finally, in Section V, we conclude with critical discussion and propose future directions.
@@ II Significance of the ‘Decoherence Functional’
Sources Arises from Noise mechanisms S($\omega$)
Charge Noise Charge fluctuators present in the defects or charge traps that reside in interfacial dielectrics, the junction tunnel barrier, and in the substrate itself. Dominant: Longitudinal ($T_{1}$), Sometimes low frequency ($T_{\phi}$) At low frequency $\propto 1/\omega$, at high frequency $\propto\omega$
Magnetic Flux Noise Stochastic flipping of qubits that reside on the surfaces of the superconductors, resulting in random fluctuations of effective-$\vec{B}$ that biases flux-tunable qubits. Longitudinal ($T_{1}$), Transverse ($T_{\phi}$) $\propto 1/\omega$
Photon Number Fluctuations In resonator, residual microwave fields in the cavity have photon number fluctuations. Transverse ($T_{\phi}$) $\propto\frac{k}{\omega^{2}+k^{2}};k:$ resonator decay rate
Quasiparticles Unpaired electrons tunneling through a qubit junction. Longitudinal ($T_{1}$), Transverse ($T_{\phi}$) $\propto\frac{N_{qp}\tau_{qp}}{\omega^{2}+\tau_{qp}^{2}};(N_{qp},\tau_{qp}):$ quasiparticle (population, lifetime)
Local Two-Level Systems (TLS) Electric dipole moments resonantly absorb energy from the oscillating $\vec{E}$ of the qubit mode, and efficiently dissipate it into the phonon or quasiparticle bath. Longitudinal ($T_{1}$), Transverse ($T_{\phi}$) Low freq.: White noise, $1/\omega$; High freq.: Lorentzian type
Table 1: Summary of common noise sources in transmon systems Krantz et al. (2019), their frequency dependencies, and noise mechanisms.
@@
$\text{Longitudinal relaxation rate:}\,\Gamma_{1}=\Gamma_{1\uparrow}+\Gamma_{1\downarrow}\equiv\frac{1}{T_{1}}$
$\text{Transversal relaxation rate:}\,\Gamma_{2}=\frac{\Gamma_{1}}{2}+\Gamma_{\phi}\equiv\frac{1}{T_{2}},$
Here, the $T_{1}$ time corresponds to the energy relaxation process, while $T_{\phi}$ represents pure dephasing, $\Gamma_{1\uparrow}$ ($\Gamma_{1\downarrow}$) denotes excitation rate (damping rate) and $\Gamma_{\phi}=1/T_{\phi}$ is the pure dephasing rate. In transmon systems, $\Gamma_{1\uparrow}\rightarrow 0$, as they are operated at millikelvin regimes. The total decoherence time $T_{2}$ reflects the loss of coherence caused by both energy relaxation and pure dephasing. The noisy density matrix $\rho_{\rm noisy}$ for the qubit after the impact of noise can then be written as Nielsen and Chuang (2011),
$\rho=\begin{pmatrix}|\alpha|^{2}&\alpha\beta^{*}\\ \alpha^{*}\beta&|\beta|^{2}\end{pmatrix}$ $\xrightarrow{\rm noise}\rho_{\rm noisy}=$
$\begin{pmatrix}1+(|\alpha|^{2}-1)e^{-\Gamma_{1}t}&\alpha\beta^{*}e^{-\Gamma_{2}t}\\ \alpha^{*}\beta e^{-\Gamma_{2}t}&|\beta|^{2}e^{-\Gamma_{1}t}\end{pmatrix}$ (1)
@@
For $1/f$ noise spectra, under free evolution, the phase decay function is itself a Gaussian $\exp[-(t/T_{\phi,G})^{2}]$ (stretched exponential decay) Krantz et al. (2019). Furthermore, this function is separable from the $T_{1}$-type exponential decay because the $T_{1}$-noise remains regular at the qubit frequency. Therefore, the modified density matrix after the action of noise is now given by Krantz et al. (2019),
@@
Here, $C(t)$ denotes the coherence curve and $F(\omega t)$ is the filter function associated with a given pulse sequence Cywiński et al. (2008). In addition to $1/f$-type dephasing mechanisms, there are also “white” pure dephasing mechanisms, which give rise to an exponential decay function for the dephasing component of $T_{2}$. The noise spectrum $S(\omega)$ thus exclusively characterizes the dephasing mechanism.
It is important to note here that the noisy density matrix in Eq. (II) is equivalent to what one may obtain using the Kraus operator-sum description of noise in quantum systems Nielsen and Chuang (2011). The Kraus operators describe the decoherence resulting from the interaction between the system and its environment over a period of time. The form of the Kraus operators takes into account the decoherence functional, as explained in the Supplementary Information (SI). The density matrix in Eq. (II) is thus identical to the density matrix in SI Eq.(3) obtained due to the combined effect of the dephasing and amplitude-damping channels on the initial density matrix, assuming that the dephasing probability $p(t)$ is simply given by $p=\frac{1}{2}(1+e^{-\Gamma_{\phi}t})$.
@@
In this case, the noisy density matrix in SI Eq.(3) gets modified to Eq. (2). This elucidates the connection between the two different approaches by demonstrating how the decoherence functional can be utilized to derive the explicit probabilities within the Kraus operator framework.
This connection makes a more precise simulation of noisy dynamics possible, bridging the gap between theoretical descriptions and practical implementations of quantum systems under the influence of noise. Therefore, critical information on qubit dynamics can be derived from the noise spectral density, $S(\omega)$, which characterizes the environment. The traditional approaches to noise spectroscopy require deconvolving the measured coherence decay $C(t)$ with the known filter function $F(\omega t)$. In other words, one has to solve an integral equation that is mathematically ill-posed and, thus, does not guarantee an accurate and unique solution. Nonetheless, this is a typical inverse problem where $C(t)$ can be readily obtained from a known $S(\omega)$. Artificial neural networks excel at solving problems such as these Wise et al. (2021); Meneses et al. (2022); Mohapatra et al. (2025), where they can find a well-approximated function to perform the deconvolution, as explained in the subsequent section. The accuracy and the uniqueness of the solution in this case rely on the constraints placed on the $S(\omega)$, however, a large parameter space can still be incorporated to retain the generality.
@@ III Noise Spectroscopy & Mitigation Methodology
The proposed methodology is designed specifically to extract the noise spectrum $S(\omega)$ associated with $T_{\phi}$ relaxation noise. Experimentally, the $T_{\phi}$ decay function is isolated by removing the $T_{1}$ contribution from the overall $T_{2}$ decay. While $T_{1}$ relaxation can be addressed primarily through improved qubit fabrication and noise-adapted QEC techniques Jayashankar and Mandayam (2023); Ofek et al. (2016), our focus here is on estimating and mitigating dephasing noise.
@@
    Training Data Generation: Based on the previous work Krantz et al. (2019), we assumed that the noise spectra $S(\omega)$ follow some complex yet well-defined functional form. Particularly, we considered that white noise dominates at relatively low frequencies, $1/\omega$-type noise becomes prominent at intermediate frequencies, and $k/(k^{2}+\omega^{2})$-type noise takes over at relatively high frequencies. Thus, we generated tens of thousands of noise spectral densities $S(\omega)$ by varying noise amplitudes and frequency cutoffs associated with the aforementioned noise types. Additionally, the stitched noise spectra were smoothed to avoid unrealistic discontinuities. We then used Eq. (3) to evaluate the corresponding decoherence curves $C(t)$ by convoluting the generated noise spectra with the CPMG filter function having 32 $\pi$-pulses (CPMG-32). Finally, we add random Gaussian noise to the synthetic coherence curves $C(t)$ to emulate the cumulative impact of experimental imperfections, collectively referred to as experimental noise. More details on data generation can be found in the SI.
@@
    Network Construction and Training: Using the numerically generated noise spectral densities and their associated coherence curves $C(t)$, we trained a convolution neural network Meneses et al. (2022) to identify the noise spectral density $S(\omega)$ based on a single decoherence curve $C(t)$ provided at the network input. A detailed description is provided in the SI.
@@
    Experimental Data Acquisition: For experimental validation, we performed measurements on IBM’s superconducting quantum processors using the Qiskit package. After determining the optimal amplitude for the square-shaped $\pi$-pulse, we performed sequential $T_{1}$ and $T_{2}$ decay experiments using a CPMG-32 pulse protocol. The choice of CPMG-32 sequence over conventional Hahn-echo protocol was made to enable probing of the complex high-frequency region of the noise spectrum, which lies beyond the white noise-dominated low-frequency region. Using the $T_{1}$ and $T_{2}$ data we generated the pure dephasing $T_{\phi}$ decay as per Eq. (5), where $P_{0}$ and $P_{1}$ are the probabilities of $|0\rangle$ (for $T_{2}$) and $|1\rangle$ (for $T_{1}$):